OpenStudy (anonymous):

how do i set this up in system of equation? A bank teller has 35 bills in $5,$10 and $20 denominations. The toatal value from these bills is $370. The total number of $10 bills and $20 bills is 7 more than the number is $5. How many bills of each denomination does the teller have?

OpenStudy (anonymous):

@UnkleRhaukus any ideas?

OpenStudy (unklerhaukus):

a+b+c=35 5a+10b+20c= 370 b+c=a+7

OpenStudy (anonymous):

okay thanks